Let X1 and X2 have the joint pmf p(x1,x2)=x1x2/36 ,x1=1,2,3 and x2=1,2,3, zero elsewhere.

Obtain the joint pmf of Y1=X1X2 and Y2=X2, and the marginal pmf of Y1.


1
Expert's answer
2021-10-15T11:45:52-0400


joint pmf of Y1=X1X2 and Y2=X2:

"p(1,1)=1\/36, p(2,2)=2\/36, p(3,3)=3\/36, p(2,1)=2\/36, p(4,2)=4\/36,"

"p(6,3)=6\/36, p(3,1)=3\/36, p(6,2)=6\/36, p(9,3)=9\/36"


marginal pmf of Y1:

"p(1)=1\/36"

"p(2)=2\/36+2\/36=4\/36"

"p(3)=3\/36+3\/36=6\/36"

"p(4)=4\/36"

"p(6)=6\/36+6\/36=12\/36"

"p(9)=9\/36"